OpenStudy (anonymous):

prove the identity (cosx/1-tanx) + (sinx/1-cotx) = cosx + sinx

OpenStudy (aravindg):

hello there! Welcome to Openstudy !

OpenStudy (anonymous):

Hi! Thanks

OpenStudy (aravindg):

ok to your question: try to write tan x and cos in terms of sin x and cos x then simplify

OpenStudy (aravindg):

*tan x and cot x

OpenStudy (anonymous):

I have it to ((cosx)/((cosx-sinx)/(cosx))+(sinx)/((sinx-cosx)/sinx))

OpenStudy (abb0t):

\[\tan(x) = \frac{ \sin(x) }{ \cos(x) }, \cot(x) = \frac{ \cos(x) }{ \sin(x) }\]

OpenStudy (aravindg):